OpenStudy (anonymous):

Leslie35: find the simple interest..principle $300 rate of interest 11% time 1 year....the simple interest is $?...please help

OpenStudy (dumbcow):

300(1+rt) = 300 + 300*rt the simple interest part is 300*rt r = rate t = time

OpenStudy (anonymous):

i put 300 it was wrong?

OpenStudy (dumbcow):

?? 300*.11*1

OpenStudy (dumbcow):

300*.11 = 33

OpenStudy (anonymous):

ty